A metalworker has a metal alloy that is 20​% copper and another alloy that is 70​% copper. How many kilograms of each alloy should the metalworker combine to create 80kg of a 60​% copper​ alloy
Found 2 solutions by Solver92311, greenestamps:
Answer by Solver92311(821)





Solve for and then calculate


John

My calculator said it, I believe it, that settles it

From
I > Ø

Answer by greenestamps(13200)



If a formal algebraic solution is not required, try this quick and easy method for solving 2-part "mixture" problems like this:

You are mixing 20% copper with 70% copper to get 60% copper.
60% is 40/50 = 4/5 of the way from 20% to 70%. (picture the three percentages on a number line, if it helps....)
That means 4/5 of the mixture is the higher percentage alloy.

ANSWER: 4/5 of 80kg, or 64kg, of 70% copper; the other 16kg of 20% copper.

CHECK:
.70(64)+.20(16) = 44.8+3.2 = 48
.60(80) = 48